OpenStudy (anonymous):

how would you be able to factor this equation? 3x^6-5x^3+2x^2

OpenStudy (anonymous):

what equation? wheres the equal sign?

OpenStudy (anonymous):

its factoring, there shouldnt be any equal sign

OpenStudy (anonymous):

take the bit that is common to all three terms "out" (of the expression)

OpenStudy (anonymous):

oh... i thought u said "factor this equation"

OpenStudy (anonymous):

that is kinda confusing..

OpenStudy (anonymous):

is there a sipliler way to explain this?O.O

OpenStudy (anonymous):

simplier*

OpenStudy (campbell_st):

do you know what a factor is..?

OpenStudy (anonymous):

I thought that was a pretty decent explanation myself....

OpenStudy (campbell_st):

thanks estudier

OpenStudy (anonymous):

umm, yep. but usually i only factor trinomials so this is new for me..

OpenStudy (campbell_st):

ok... can you simplify \[x^4 \times x^2 = \]

OpenStudy (anonymous):

Um...it is a trinomial isn't it?

OpenStudy (campbell_st):

yep a 3 term expression

OpenStudy (anonymous):

2x^6?

OpenStudy (campbell_st):

nope... try again

OpenStudy (anonymous):

lol i cant . tell me the answer xD

OpenStudy (campbell_st):

ok... here is an index law you should know... \[x^a \times x^b = x^{a+b}\] now back to the question \[x^4 \times x^2 =\]

OpenStudy (anonymous):

x^6?

OpenStudy (campbell_st):

correct.. so the 1st term in your expression is \[3\times x^6 = 3 \times x^4 \times x^2\] the 2nd term is \[-5 \times x^3 = -5 \times x \times x^2\] the 3rd term is \[2 \times x^2 \] in the expanded form above... what is the common factor..?

OpenStudy (anonymous):

2(x)(x)?

OpenStudy (campbell_st):

no... then is only 1 thing that is contained in the 3 terms written in expanded form... can you identify it..

OpenStudy (anonymous):

no idea...

OpenStudy (campbell_st):

\[x^2\] is common to all 3 terms... so that is the common factor \[3x^4\times x^2 - 5x \times x^2 + 2\times x^2 = x^2(3?? -5?? + 2)\] so what goes inside the brackets with 3 and - 5

OpenStudy (anonymous):

ummm x^4 and x?

OpenStudy (campbell_st):

correct so the simplified form is \[x^2(3x^4 - 5x + 2)\]

OpenStudy (anonymous):

yay! i get it now !

OpenStudy (anonymous):

thankyou

OpenStudy (campbell_st):

good luck